A mobile company charges Rs 1.5 per minute during the daytime (7 am to 11 pm) and Rs x per minute during night (11 pm to 7 am). The total talking time during night decreases by 50%. Find the value of x if the average amount charged by the mobile company during the whole day is Rs 1.40 per minute?
1
Expert's answer
2019-10-16T11:02:04-0400

Let the talk time during daytime be "y" minutes.

As the  total talking time during night decreases by 50% with respect to daytime.So,it will be half of that in daytime.

Talking time for 11 pm to 7 am will be "\frac{y}{2}" minutes.

Total talking time"=y+\frac{y}{2}=\frac{3y}{2}"

As per the question,Total amount of charge during daytime "=y\times 1.5"

Total amount of charge during nighttime"=x\times \frac{y}{2}"

Average amount charged"=\frac{Total \ amount}{Total \ time}=\frac{y \times 1.5+x\times 0.5y}{1.5y}=1.40"

"\implies \frac{1.5+0.5x}{1.5}=1.40 \implies x=Rs\ 1.20/min"